OpenStudy (anonymous):

What is the inverse of the relation described by y = (2x - 3y)^2?

OpenStudy (anonymous):

oops, typed it wrong y = (2x - 3)^2

OpenStudy (anonymous):

switch x and y, then solve for y.. x=(2y-3)^2 wanna try and start?

OpenStudy (anonymous):

ok x = (2y-3)^2 sqrt(x) = 2y - 3 so far so good?

OpenStudy (anonymous):

Yep, keep going

OpenStudy (anonymous):

sqrt(x) = 2y - 3 sqrt(x) + 3 = 2y [sqrt(x) + 3]/2 = y

OpenStudy (anonymous):

is that my answer?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome